OpenStudy (anonymous):

If C(x) = (x - 2)^2 + 5, find the maximum profit (which might be negative!). The maximum profit is

OpenStudy (freckles):

So what is the function for P if we know C?

OpenStudy (anonymous):

you can see from this form of the equation that the vertex is at (2,5). 5 is the min profit because the vertex is at the minimum. There is no max profit because each side of the parabola tends toward infinity when x gets larger positively or negatively

OpenStudy (anonymous):

Oh this is the cost function, not the profit function

OpenStudy (amistre64):

x*C(x) = profit right?

OpenStudy (freckles):

P(x)=R(x)-C(x)=x p(x)-C(x) but what is the little p guy?

OpenStudy (freckles):

little p is the demand function

OpenStudy (freckles):

so i guess C is all we know?